a pharmacist needs 500 milligrams of a 20% phenobarbital solution but has only 5% and 25% phenobarbital solutions available. find out how many milligrams of each she should mix to get the desired solution?
Answer by stanbon(75887)

a pharmacist needs 500 milligrams of a 20% phenobarbital solution but has only 5% and 25% phenobarbital solutions available. find out how many milligrams of each she should mix to get the desired solution?
-------
Equation:
pheno + pheno = pheno
0.05x + 0.25(500-x) = 0.20*500
-----
5x + 25*500 - 25x = 20*500
-20x = -5*500
----
x = 500/4
x = 125 ml (amt of 5% solution needed)
500-x = 375 ml (amt of 25% solution needed)
-----------
Cheers,
Stan H.
-------